# GS-Occ3D: Scaling Vision-only Occupancy Reconstruction for Autonomous Driving with Gaussian Splatting

### v1

@@ Abstract
Occupancy is crucial for autonomous driving, providing essential geometric priors for perception and planning. However, existing methods predominantly rely on LiDAR-based occupancy annotations, which limits scalability and prevents leveraging vast amounts of potential crowdsourced data for auto-labeling. To address this, we propose GS-Occ3D, a scalable vision-only framework that directly reconstructs occupancy. Vision-only occupancy reconstruction poses significant challenges due to sparse viewpoints, dynamic scene elements, severe occlusions, and long-horizon motion. Existing vision-based methods primarily rely on mesh representation, which suffer from incomplete geometry and additional post-processing, limiting scalability. To overcome these issues, GS-Occ3D optimizes an explicit occupancy representation using an Octree-based Gaussian Surfel formulation, ensuring efficiency and scalability. Additionally, we decompose scenes into static background, ground, and dynamic objects, enabling tailored modeling strategies: (1) Ground is explicitly reconstructed as a dominant structural element, significantly improving large-area consistency; (2) Dynamic vehicles are separately modeled to better capture motion-related occupancy patterns. Extensive experiments on the Waymo dataset demonstrate that GS-Occ3D achieves state-of-the-art geometry reconstruction results. By curating vision-only binary occupancy labels from diverse urban scenes, we show their effectiveness for downstream occupancy models on Occ3D-Waymo and superior zero-shot generalization on Occ3D-nuScenes. It highlights the potential of large-scale vision-based occupancy reconstruction as a new paradigm for autonomous driving perception. Project Page.
@@ 1 Introduction
Goal: Vision-only reconstruction. Current occupancy reconstruction methods for autonomous driving primarily depend on LiDAR-based annotations [30, 104]. As shown in Fig.˜1, these approaches face inherent scalability limitations due to the high costs associated with data collection and annotation. In contrast, vision-centric frameworks offer a promising alternative by leveraging large-scale crowdsourced data for self-supervised auto-labeling, eliminating the reliance on expensive LiDAR sensors [71, 54, 81, 41, 47]. Despite its potential, vision-only occupancy reconstruction remains a challenging task due to the sparsity of viewpoints, severe occlusions, dynamic scene elements, and long-horizon motion. For example, a typical scene in the Waymo dataset contains approximately 1,000 frames captured from only five cameras, resulting in limited co-visible regions across frames and incomplete scene observations.
@@
Limitations of prior methods. (1) Prior works focusing on geometry reconstruction [9, 16, 21, 96, 27] primarily target indoor, object-centric, or simple outdoor scenes. Extending these techniques to large-scale driving scenes leads to substantial accuracy degradation, particularly in weakly textured regions and along long-horizon, high-speed trajectories. (2) Existing methods that attempt street scene geometry reconstruction [64, 51, 22, 15] typically rely on mesh-based representations, which introduce overly smooth surfaces (in implicit methods) or fragmented, hole-ridden reconstruction (in GS-based methods). These issues necessitate extensive post-processing, hindering their scalability for efficient vision-only label curation. (3) Last but not least, most prior approaches [51, 22, 64, 99, 27, 75, 95, 27, 96] also focus solely on static scenes, failing to capture the dynamic aspects of real-world driving environments. This lack of dynamic object modeling reduces their applicability to downstream autonomous driving tasks [98, 67, 46, 105].
To summarize, scalable vision-only occupancy reconstruction for autonomous driving is constrained by three major challenges: (1) Limited geometric priors from sparse viewpoints, making accurate 3D reconstruction difficult. (2) Degradation over long trajectories, leading to inconsistent and incomplete geometry. (3) Dynamic occlusions from moving objects, creating challenges in accurately modeling interactions in complex urban environments.
@@
By utilizing this tailored representation, GS-Occ3D ensures high spatial-temporal consistency and preserves multi-scale geometric fidelity across long-horizon sequences. With this approach, we reconstruct the entire Waymo dataset to generate vision-only point clouds. These point clouds are then processed through a pipeline consisting of frame-wise division, multi-frame aggregation, and voxelization. This enables the curation of vision-only occupancy labels, facilitating the training of state-of-the-art downstream occupancy models.
@@ 2 Related Works
For dynamic scenes, SUDS [74] and EmerNeRF [88] model all elements within a single field, while some methods [58, 91, 84, 72] decompose scenes into foreground and background for better motion handling. Recent GS-based works [87, 108, 107, 106, 11, 25, 65, 38, 77, 85] further improve the fidelity and efficiency. More recently, some works [36, 53, 93, 23, 90, 57] integrate diffusion models [26, 17] for diverse driving scene generation. However, these methods rely on LiDAR or prioritize rendering over geometry. We address this gap with a vision-only reconstruction method focused on geometry.
Vision-only Geometry Reconstruction. Vision-only geometry reconstruction in urban street scenes is challenging due to dynamic objects, sparse views, occlusion, and long-horizon motion. NeRF-based methods like DNMP [51] use neural mesh primitives, while StreetSurf [22] incorporates monocular cues (e.g., depth, normal) from pretrained models. However, their reliance on complete meshes or full-volume processing limits scalability. GS-based methods have advanced object-centric and bounded scene reconstruction [9, 16, 21, 96, 99, 27, 75, 95], but scaling to large street scenes introduces issues like uneven point distribution, holes, and floaters, making them unsuitable for downstream tasks. Recent grid-based improvements [64, 15] enhance mesh quality but struggle in weak-texture regions. In contrast, our approach reconstructs both ground and dynamic geometry, advancing scalable vision-only occupancy reconstruction and enabling seamless label curation for downstream autonomous driving tasks.
@@
Most methods rely on costly LiDAR-based occupancy labels. LiDAR-based methods [66, 63, 37] complete scene occupancy from sparse LiDAR inputs, while camera-based methods predict 3D occupancy by extracting features from 2D images [6, 40, 39, 43, 102, 28, 42, 94, 59, 50, 83, 61, 12]. Monocular methods infer 3D structure from a single image via 2D-to-3D backprojection [6] or depth-aware cross-attention [40], while multi-view approaches generate 3D volumetric features from multiple camera perspectives [39, 43, 102, 28, 42, 94, 59, 50, 83]. Some methods leverage 3DGS to transform 2D images into dense gaussian representations [30, 104]. Despite strong results, these methods heavily depend on LiDAR-based annotations, which is costly and time-consuming. Recent vision-centric approaches aim to reduce reliance on LiDAR by using volume rendering with 2D supervision [29, 60, 100, 61, 103, 4, 7, 49] or pretrained vision-language models [109, 3, 31]. Instead, we focus on leveraging vision-only 3D geometry labels, which is cheaper, more scalable and efficient for downstream 3D occupancy models.
@@ 3 Methods
Fig.˜2 illustrates the overview of our framework. We first generate a sparse point cloud and ground surfels from panoramic street views captured along long trajectories. For scalable vision-only geometry reconstruction, we use an octree-based Gaussian Surfel representation integrating ground, background, and dynamic objects. Our ground-truth curation pipeline refines the vision-only point cloud through frame-wise division and multi-frame aggregation, increasing density, especially for dynamic objects. Ray-casting then resolves occlusions for accurate voxel occupancy labeling. The resulting vision-only labels can supervise downstream occupancy models, improving their generalization and geometric reasoning capabilities.
@@ 3.1 Scalable Vision-only Geometry Reconstruction
Preprocessing. Given the absence of geometric priors derived from LiDAR, our pipeline exclusively uses detector-free SfM [68, 24] and ground gaussians detailed later to establish sparse point clouds as the initial scene representation. We also use an off-the-shelf segmentation model [14] for necessary decoupling.
@@
We adopt an adaptive anchor control strategy inspired by [52, 62] to manage surfels in large scenes. To further enhance geometry completeness across scales, we apply cumulative LOD levels instead of a single LOD, allowing the representation to capture both coarse scene coverage and fine geometric details, ensuring high-fidelity reconstruction across varying spatial scales.
Ground Reconstruction. Preliminary experiments show that existing geometry reconstruction methods, despite incorporating geometric constraints, struggle with weakly-textured ground regions in sparse-view street scenes, underscoring the need for a dedicated approach to enhance ground modeling accuracy.
Assuming the road surface is approximately parallel to the camera poses, we initialize Ground Gaussian surfels by projecting camera poses onto the xy-plane, inspired by [19]. To handle elevation, each surfel’s z-coordinate is adjusted using the nearest camera pose with a fixed height offset, while its orientation inherits the nearest camera’s rotation. This initialization aligns surfels with the road slope, enabling adaptation to various terrains as shown in Fig.˜3. Finally, planar regularization encourages smooth planar structures, further improving ground geometry as shown in Fig.˜8.
Dynamic Reconstruction. We assume each dynamic vehicle is associated with a 3D bounding box predicted from RGB images and represented by a point cloud with tracked poses $\mathbf{R}_{t}$ and $\mathbf{t}_{t}$. We initialize dynamic vehicles using vision-based 3D object tracking method [20]. A fixed number of points are then sampled within each box.
@@
where $\Delta\mathbf{t}_{t}$ is a learnable translation vector, and $\Delta\mathbf{R}_{t}$ is a rotation matrix constructed from a learnable yaw offset $\Delta\theta_{t}$. This design allows direct gradient computation, ensuring efficient training.
Loss Function. The total loss is a weighted sum of five components: RGB loss, geometry loss, object loss, ground loss, and sky loss, expressed as:
@@ 3.2 Occupancy Labels Curation
With scalable geometry reconstruction, it becomes feasible to construct vision-only 3D occupancy labels. However, the relatively sparse point cloud produced during geometry reconstruction makes it difficult to obtain dense voxel representations, which leads to the sparsity problem. As the point cloud is densified, identifying occluded and invisible voxels from the camera’s viewpoint becomes essential. To address these issues, our label curation pipeline employs frame-wise division and multi-frame aggregation to define per-frame perception ranges for vision-only point clouds and increase point-cloud density, particularly for dynamic objects with incomplete observations. Then we can operate ray-casting to determine the occupancy status of every voxel, explicitly handling occlusion.
@@
Directly merging points across frames can cause smearing or distortion in dynamic objects like vehicles, so it is necessary to process them separately. Since our pipeline explicitly separates static and dynamic components, we can extract optimized points within tracked bounding boxes without additional segmentation. These points are transformed from the sensor coordinate system to the box coordinate system, following a process similar to [71]. By concatenating these transformed points across frames, we effectively densify the point clouds for dynamic objects.
@@
Voxelization. To generate a 3D occupancy grid from aggregated point clouds, a straightforward approach is to mark voxels containing points as occupied and others as free. However, due to the limited camera field of view, some occupied voxels are only partially observed or entirely occluded from the camera’s perspective, which can lead to incorrect labeling as free. This ambiguity can confuse downstream models during training, making it necessary to distinguish between free and unobserved voxels.
Inspired by [71], we utilize a ray-casting operation to determine the visibility of each voxel. Specifically, we trace a ray from the camera origin to the center of each occupied voxel. Along each ray, the first occupied voxel encountered is labeled as observed, while the rest are labeled as unobserved. Any voxel that is not traversed by any camera ray is also considered unobserved.
@@ 4 Experiments
Figure 3: Visualization of Geometry Reconstruction on Waymo. The color represents CD with respect to LiDAR, ranging from blue (lower CD) to red (higher CD). Our method exhibits improved reconstruction fidelity in weakly-textured regions compared to other methods, while maintaining structural completeness comparable to LiDAR point cloud, even in the absence of geometric priors.
@@ 4.1 Geometry Reconstruction Experimental Setup
Datasets. To evaluate performance and scalability in large-scale open scenes, we use the Waymo Open Dataset [69]. We use all five camera views and all frames in each scene, resulting in about 1,000 images per scene. LiDAR point clouds are used as reference to assess geometric accuracy.
Baselines. For geometry reconstruction, we evaluate our method against state-of-the-art implicit methods (NeuS [78], F2-NeRF [79], StreetSurf [22]) and GS-based methods (2DGS [27], PGSR [9], GVKF [64]). All explicit baselines use the same input point clouds for initialization.
@@
4.2 Geometry Reconstruction Result Analysis
SOTA Geometry Reconstruction. Footnote˜3 shows that we achieve SOTA geometry reconstruction, surpassing methods that use monocular cues or LiDAR-input or supervised like StreetSurf and NeuS, which demonstrates the reliability of our upstream geometry reconstruction. In addition to reconstruction quality, our method also maintains competitive rendering quality and training efficiency. This is partly due to the octree-based representation, which is both memory efficient and faster to train. Although our method reconstructs denser point clouds, the structured octree helps accelerate geometry convergence while ensuring high rendering and geometric accuracy. Under sparse-view conditions, the octree-based Gaussian surfel with LoD preserves scene geometric fidelity through multi-scale consistency.
Comparable and Complementary to LiDAR. As shown in  Fig.˜3 and  Fig.˜4, our vision-only point cloud is globally comparable to LiDAR scans. In certain regions, such as tall buildings and thin poles, our reconstruction achieves even higher quality, capturing finer details. Additionally, it provides more detailed textures and reconstructs areas beyond LiDAR’s limited range, making it reliable for downstream tasks. As observed in the comparison, our reasonable reconstruction beyond the LiDAR coverage is also counted as regions with high CD.
@@ 4.2 Geometry Reconstruction Result Analysis
Through reconstructing the full Waymo Open Scene dataset, we observe that vision-only methods fail to handle ego-static scenes. As a result, we exclude these unreliable regions from the labels used for training occupancy reconstruction models. The experiments compare downstream occupancy reconstruction models trained with the labels from Occ3D [71] and our produced labels. We evaluate the generalization performance on the validation set of Occ3D-Waymo and Occ3D-nuScenes and the fitting performance on the training set, aiming to assess the reliability of vision-only 3D occupancy reconstruction.
[…]
@@
Metrics. To evaluate the geometric accuracy and reliability of our labels, we use the Intersection over Union (IoU) to assess geometry performance of 3D binary occupancy prediction. For fair comparison, the evaluation only considers voxels within the visible region of the camera views.
@@
Label Curation Results.  Fig.˜5 visualizes the comparison between our labels and Occ3D. Without geometric priors and using only a sparse point cloud for initialization, our method still achieves globally comparable geometry to the LiDAR-based labels. This ensures reliable geometry supervision for training the occupancy model.
Comparable and Superior Zero-shot Occupancy Generalization Results.  Fig.˜6 and  Tab.˜2 show the generalization results of CVT-Occ [94] on the validation set. Our method demonstrates the ability to reconstruct generalized geometry, achieving reasonable and overall comparable results on Occ3D-Waymo. Despite inherent camera limitations (e.g., Waymo’s forward-facing views versus LiDAR’s 360-degree coverage), our method achieves slightly lower performance than Occ3D but still within a reasonable range. Notably, we demonstrate superior zero-shot generalization on nuScenes with diverse camera settings, yielding more complete geometry in textured and distant regions. This is particularly impressive given that Occ3D relies on high-end LiDAR data. Our method relies purely on camera inputs yet delivers comparable or even better generalization performance in certain setups. This highlights not only the scalability of vision-only methods, but also their potential to match or surpass LiDAR-based baselines in real world.
Occupancy Fitting Results. The last two rows of Tab.˜2 present the fitting results of CVT-Occ [94] on different training sets. The performance gap evaluated on Occ3D-Train and Ours-Train is small, and the precision on the Occ3D-Train and Occ3D-Val remains similar. This indicates that our labels provide learnable geometry, enabling the model to capture meaningful geometry information.
@@
Advantages over LiDAR. Based on results above, we analyze the advantages of vision-only labels compared to LiDAR labels: (1) Wider coverage: Vision-only geometry reconstruction can mitigate the limited spatial coverage of LiDAR, especially over large-scale areas. In certain cases, it even outperforms LiDAR, such as reconstructing high-rise buildings. (2) Superior zero-shot generalization: Compared to LiDAR-based labels, downstream occupancy models trained with vision-only labels generalize to a wider range of geometries and have superior zero-shot generalization ability. (3) Cheap and Rich Semantics: RGB images inherently capture color, texture, and object-class cues, unlike LiDAR semantics which require labor-intensive annotation or costly multi-sensor fusion. Leveraging rich 2D visual semantics via Mask2Former [14], we reconstruct 3D labels with a richer category set (up to 66 vs. Occ3D’s 16), including small objects and categories missed by Occ3D. Fig.˜7 shows semantics difficult for LiDAR to obtain yet vital for driving. As highlighted, unlike Occ3D, we can distinguish motorcycles and identify lane markings and crosswalks. (4) Greater potential in adverse weather: Although adverse weather impacts both sensors, vision-based systems can leverage rich textures and semantics in the image domain, along with learned priors from large-scale visual data, to reconstruct degraded scenes more effectively. From Fig.˜7, we even generalize better than LiDAR-based baseline in rainy scenes.

### v3

# GS-Occ3D: Scaling Vision-only Occupancy Reconstruction with Gaussian Splatting

@@ Abstract
Occupancy is crucial for autonomous driving, providing essential geometric priors for perception and planning. However, existing methods predominantly rely on LiDAR-based occupancy annotations, which limits scalability and prevents leveraging vast amounts of potential crowdsourced data for auto-labeling. To address this, we propose GS-Occ3D, a scalable vision-only framework that directly reconstructs occupancy. Vision-only occupancy reconstruction poses significant challenges due to sparse viewpoints, dynamic scene elements, severe occlusions, and long-horizon motion. Existing vision-based methods primarily rely on mesh representation, which suffer from incomplete geometry and additional post-processing, limiting scalability. To overcome these issues, GS-Occ3D optimizes an explicit occupancy representation using an Octree-based Gaussian Surfel formulation, ensuring efficiency and scalability. Additionally, we decompose scenes into static background, ground, and dynamic objects, enabling tailored modeling strategies: (1) Ground is explicitly reconstructed as a dominant structural element, significantly improving large-area consistency; (2) Dynamic vehicles are separately modeled to better capture motion-related occupancy patterns. Extensive experiments on the Waymo dataset demonstrate that GS-Occ3D achieves state-of-the-art geometry reconstruction results. By curating vision-only binary occupancy labels from diverse urban scenes, we show their effectiveness for downstream occupancy models on Occ3D-Waymo and superior zero-shot generalization on Occ3D-nuScenes. It highlights the potential of large-scale vision-based occupancy reconstruction as a new paradigm for scalable auto-labeling. Project Page.
@@ 1 Introduction
Goal: Vision-only reconstruction. Current occupancy reconstruction methods primarily depend on LiDAR-based annotations [30, 104]. As shown in Fig.˜1, these approaches face inherent scalability limitations due to the high costs associated with data collection and annotation. In contrast, vision-centric frameworks offer a promising alternative by leveraging large-scale crowdsourced data for self-supervised auto-labeling, eliminating the reliance on expensive LiDAR sensors [71, 54, 81, 41, 47]. Despite its potential, vision-only occupancy reconstruction remains a challenging task due to the sparsity of viewpoints, severe occlusions, dynamic scene elements, and long-horizon motion. For example, a typical scene in the Waymo dataset contains approximately 1,000 frames captured from only five cameras, resulting in limited co-visible regions across frames and incomplete scene observations.
@@
Limitations of prior methods. (1) Prior works focusing on geometry reconstruction [9, 16, 21, 96, 27] primarily target indoor, object-centric, or simple outdoor scenes. Extending these techniques to large-scale driving scenes leads to substantial accuracy degradation, particularly in weakly textured regions and along long-horizon, high-speed trajectories. (2) Existing methods that attempt street scene geometry reconstruction [64, 51, 22, 15] typically rely on mesh-based representations, which introduce overly smooth surfaces (in implicit methods) or fragmented, hole-ridden reconstruction (in GS-based methods). These issues necessitate extensive post-processing, hindering their scalability for efficient vision-only label curation. (3) Last but not least, most prior approaches [51, 22, 64, 99, 27, 75, 95, 27, 96] also focus solely on static scenes, failing to capture the dynamic aspects of real-world driving environments. This lack of dynamic object modeling reduces their applicability to downstream tasks [98, 67, 46, 105].
To summarize, scalable vision-only occupancy reconstruction is constrained by three major challenges: (1) Limited geometric priors from sparse views, making accurate reconstruction hard. (2) Degradation over long trajectories, leading to inconsistent and incomplete geometry. (3) Dynamic occlusions from moving objects, creating challenges in accurately modeling interactions in complex scenes.
@@
By utilizing this tailored representation, we ensure high spatial-temporal consistency and preserve multi-scale geometric fidelity across long-horizon sequences. With this approach, we reconstruct the entire Waymo dataset to generate vision-only point clouds, which are then processed through a pipeline consisting of frame-wise division, multi-frame aggregation and voxelization. This enables the curation of vision-only occupancy labels, facilitating the training of state-of-the-art downstream occupancy models.
@@ 2 Related Works
For dynamic scenes, some works [74, 88] model all elements within a single field, while some works [58, 91, 84, 72] decompose scenes into foreground and background for better motion handling. Recent GS-based works [87, 108, 107, 106, 11, 25, 65, 38, 77, 85] further improve the fidelity and efficiency. More recently, some works [36, 53, 93, 23, 90, 57] integrate diffusion models [26, 17] for diverse driving scene generation. However, they rely on LiDAR or prioritize rendering over geometry. We address this gap with a vision-only reconstruction method focused on geometry.
Vision-only Geometry Reconstruction. Vision-only geometry reconstruction in urban scenes is challenging due to dynamic objects, sparse views, occlusion and long-horizon motion. NeRF-based methods like DNMP [51] use neural mesh primitives, while StreetSurf [22] incorporates monocular cues (e.g., depth, normal) from pretrained models. However, their reliance on complete meshes or full-volume processing limits scalability. GS-based methods advance object-centric and bounded scene reconstruction [9, 16, 21, 96, 99, 27, 75, 95], but scaling to street scenes faces issues like uneven point distribution, holes, and floaters, making them unsuitable for downstream tasks. Recent grid-based improvements [64, 15] enhance mesh quality but struggle in weak-texture regions. In contrast, we reconstruct both ground and dynamic geometry, advancing scalable vision-only occupancy reconstruction and enabling seamless label curation for downstream tasks.
@@
Most methods rely on costly LiDAR-based occupancy labels. LiDAR-based methods [66, 63, 37] complete scene occupancy from sparse LiDAR inputs, while camera-based methods predict 3D occupancy by extracting features from 2D images [6, 40, 39, 43, 102, 28, 42, 94, 59, 50, 83, 61, 12]. Monocular methods infer 3D structure from a single image via 2D-to-3D backprojection [6] or depth-aware cross-attention [40], while multi-view approaches generate 3D volumetric features from multiple camera views [39, 43, 102, 28, 42, 94, 59, 50, 83]. Some works use GS to transform 2D images into dense gaussian representations [30, 104]. Despite strong results, they heavily depend on LiDAR-based annotations, which is costly and time-consuming. Recent vision approaches reduce reliance on LiDAR by using volume rendering with 2D supervision [29, 60, 100, 61, 103, 4, 7, 49] or pretrained vision-language models [109, 3, 31]. Instead, we focus on leveraging vision-only 3D geometry labels, which is cheaper, more scalable and efficient for downstream occupancy models.
@@ 3 Methods
Fig.˜2 illustrates the overview of our framework. We first generate a sparse point cloud and ground surfels from panoramic street views captured along long trajectories. For scalable vision-only geometry reconstruction, we use an octree-based Gaussian Surfel representation integrating ground, background, and dynamic objects. Our label curation pipeline refines the vision-only point cloud through frame-wise division and multi-frame aggregation, increasing density, especially for dynamic objects. Ray-casting then resolves occlusions for accurate voxel occupancy labeling. The resulting vision-only labels can supervise downstream occupancy models, improving their generalization and geometric reasoning capabilities.
@@ 3.1 Scalable Vision-only Geometry Reconstruction
Preprocessing. In the absence of geometric priors, we exclusively use detector-free SfM [68, 24] and ground gaussians detailed later to generate sparse point clouds as the initial scene representation. We also use an off-the-shelf segmentation model [14] for necessary decoupling.
@@
We use an adaptive anchor control strategy inspired by [52, 62] to manage surfels. To enhance geometry completeness across scales, we apply cumulative LOD levels instead of a single one, to capture both coarse scene coverage and fine geometric details for high-fidelity reconstruction.
Ground Reconstruction. Existing reconstruction methods, despite incorporating geometric constraints, struggle with weakly-textured ground in sparse-view scenes, highlighting the need for a dedicated ground modeling approach.
Assuming the road is parallel to the cameras, we initialize Ground Gaussian surfels by projecting camera poses onto the xy-plane, inspired by [19]. To handle elevation, each surfel’s z-coordinate is adjusted using the nearest camera pose with a fixed height offset, while its orientation inherits the nearest camera’s rotation. This initialization aligns surfels with the road slope, enabling adaptation to various terrains as shown in Fig.˜3. Finally, planar regularization encourages smooth planar structures, further improving ground geometry as shown in Fig.˜8.
Dynamic Reconstruction. We assume each dynamic vehicle is associated with a 3D bounding box predicted from RGB images and represented by points with tracked poses $\mathbf{R}_{t}$ and $\mathbf{t}_{t}$. We initialize dynamic vehicles using vision-based 3D object tracking method [20]. A fixed number of points are then sampled within each box.
@@
The learnable translation $\Delta\mathbf{t}_{t}$ and rotation $\Delta\mathbf{R}_{t}$ enable efficient training via direct gradient computation.
Loss Function. The total loss combines weighted RGB, geometry, object, ground, and sky components:
@@ 3.2 Occupancy Labels Curation
With scalable geometry reconstruction, it becomes feasible to curate vision-only 3D occupancy labels. However, the relatively sparse point cloud produced during geometry reconstruction makes it hard to obtain dense voxel representations, which leads to the sparsity problem. As the point cloud is densified, identifying occluded and invisible voxels from the camera views becomes essential. To address these issues, we employ frame-wise division and multi-frame aggregation to define per-frame perception ranges for vision-only point clouds and increase point-cloud density, particularly for dynamic objects with incomplete observations. Then we can operate ray-casting to determine the occupancy status of every voxel, explicitly handling occlusion.
@@
Directly merging points across frames can cause smearing or distortion in dynamic objects like vehicles, so it is necessary to process them separately. Since our pipeline explicitly decouples static and dynamic components, we can extract optimized points within tracked bounding boxes without additional segmentation. These points are transformed from the sensor coordinate system to the box coordinate system, following a process similar to [71]. By concatenating these transformed points across frames, we effectively densify the point clouds for dynamic objects.
@@
Voxelization. To generate a 3D occupancy grid from aggregated point clouds, a straightforward approach is to mark voxels containing points as occupied and others as free. However, due to the limited camera field of view, some occupied voxels are only partially observed or entirely occluded from the camera’s perspective, which can lead to incorrect labeling as free. This ambiguity can confuse downstream models during training, making it necessary to distinguish between free and unobserved voxels. Inspired by [71], we use ray-casting to determine voxel visibility. We cast a ray from the camera to each occupied voxel. Only the first occupied voxel encountered along any ray is labeled as observed, while all others are considered unobserved.
@@ 4 Experiments
Figure 3: Visualization of Geometry Reconstruction on Waymo. The color represents CD with respect to LiDAR, ranging from blue (lower CD) to red (higher CD). We exhibit improved reconstruction fidelity in weakly-textured regions compared to other methods, while maintaining structural completeness comparable to LiDAR point cloud, even in the absence of geometric priors.
@@ 4.1 Geometry Reconstruction Experimental Setup
Datasets. To evaluate performance and scalability in large-scale open scenes, we use the Waymo dataset [69]. We use all five camera views and all frames in each scene, resulting in about 1,000 images per scene. LiDAR point clouds are used as reference to assess geometric accuracy.
Baselines. For geometry reconstruction, we evaluate our method against SOTA implicit methods (NeuS [78], F2-NeRF [79], StreetSurf [22]) and GS-based methods (2DGS [27], PGSR [9], GVKF [64]). All explicit baselines use the same input point clouds for initialization.
@@
4.2 Geometry Reconstruction Result Analysis
SOTA Geometry Reconstruction. Footnote˜3 shows that we achieve SOTA geometry reconstruction, surpassing methods that use monocular cues or LiDAR-input or supervised like StreetSurf and NeuS, which demonstrates the reliability of geometry reconstruction. Besides geometry, we also maintain competitive rendering quality and training efficiency. This is partly due to the octree-based representation, which is both memory efficient and faster to train. Although we reconstruct denser point clouds, the structured octree helps accelerate geometry convergence while ensuring high rendering and geometric accuracy. Under sparse-view conditions, the octree-based gaussian surfel with LoD preserves geometric fidelity through multi-scale consistency.
Comparable and Complementary to LiDAR. As shown in  Fig.˜3 and  Fig.˜4, our point cloud is globally comparable to LiDAR. In certain regions, such as tall buildings and thin poles, we achieve even higher quality, capturing finer details. Additionally, it provides more detailed textures and reconstructs areas beyond LiDAR’s limited range, making it reliable for downstream tasks. As observed in the comparison, our reasonable reconstruction beyond the LiDAR coverage is also counted as regions with high CD.
@@ 4.2 Geometry Reconstruction Result Analysis
Through reconstructing the full Waymo dataset, we find vision-only methods fail to handle ego-static scenes. So we exclude these unreliable parts from the labels for training occupancy models. The experiments compare downstream occupancy models trained with the labels from Occ3D [71] and our labels. We evaluate the generalization performance on the validation set of Occ3D-Waymo and Occ3D-nuScenes and the fitting performance on the training set.
Datasets. Occ3D-Waymo is among the most diverse and comprehensively labeled open-source 3D occupancy datasets [71], which contains 798 training scenes and 202 validation scenes, covering approximately 200,000 frames in total. For fairness, we exclude ego-static scenes that cannot be reliably handled by vision-only reconstruction, resulting in 637 training scenes and 165 validation scenes. We also use original Occ3D-nuScenes validation set including 150 scenes for evaluating zero-shot generalization.
@@
Metrics. To evaluate the geometric accuracy and reliability of our labels, we use the Intersection over Union (IoU) to assess geometry performance of 3D occupancy prediction. For fair comparison, the evaluation only considers voxels within the visible region of the camera views.
@@
Label Curation Results.  Fig.˜5 visualizes the comparison between our labels and Occ3D. Without geometric priors and initializing with sparse point clouds, we still achieve globally comparable geometry to the LiDAR-based labels. This ensures reliable geometry for downstream models.
Comparable and Superior Zero-shot Occupancy Generalization Results.  Fig.˜6 and  Tab.˜2 show the generalization results of CVT-Occ [94] on the validation set. We demonstrate the ability to reconstruct generalizable geometry, achieving reasonable and overall comparable results on Occ3D-Waymo. Despite inherent camera limitations (e.g., Waymo’s forward-facing views versus LiDAR’s 360-degree coverage), we achieve slightly lower performance on the Occ3D-Val(Waymo) but still within a reasonable range. While both we and Occ3D perform better on our respective validation sets, we notably demonstrate superior zero-shot generalization on nuScenes with diverse camera settings, yielding more complete geometry in textured and distant regions. This is particularly impressive given that Occ3D relies on high-end LiDAR data. We rely purely on camera inputs yet delivers comparable or even better generalization performance in certain setups. This highlights not only the scalability of vision-only methods, but also their potential to match or surpass LiDAR-based baselines in real world.
Occupancy Fitting Results. As shown in the last two rows of Tab.˜2, the performance gap evaluated on Occ3D-Train and Ours-Train is small, and the precision on the Occ3D-Train and Occ3D-Val remains similar. This indicates that our labels provide learnable geometry, enabling the model to capture meaningful geometry information.
@@
Advantages over LiDAR. (1) Wider coverage: Vision-only reconstruction mitigates LiDAR’s limited spatial range, excelling in large-scale areas and tall structures like high-rise buildings. (2) Superior zero-shot generalization: Models trained with our labels can predict a wider range of geometries and generalize better to unseen scenes than LiDAR-based ones. (3) Cheap and Rich Semantics: Unlike LiDAR semantics which require labor-intensive annotation or costly multi-sensor fusion, RGB images inherently capture color, texture, and object-class cues. Leveraging 2D segmentation models like  [14], we produce a richer 3D category set (66 vs. Occ3D’s 16), identifying objects LiDAR misses yet vital for driving like motorcycles, lane markings and crosswalks. (Fig.˜7). (4) Greater potential in adverse weather: Although adverse weather impacts both sensors, vision can use rich textures and semantics and learned priors from large-scale data, to reconstruct degraded scenes more effectively. From Fig.˜7, we even generalize better than LiDAR in rainy scenes.